OpenStudy (anonymous):

Become a fan You're a fanNot a fan The perimeter of a rectangle is 110m , and the area of the rectangle is 54m^2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

im pretty posivite that u do base x Hight

OpenStudy (anonymous):

2x+2y=110 54=x*y y=54/x 110=2(54/x)+2x 55=(54/x)+x 55x-x^2)=54 -1x^2+55x-54=0 use quadratic to find value for x=1 or 54 1 obviously can't be the answer so plug 54 back in for x to find y

OpenStudy (anonymous):

that can't be right you cant have an area half the value of your perimeter